OpenStudy (anonymous):

23780 LB. how much in t.

OpenStudy (lgbasallote):

huh?

OpenStudy (lgbasallote):

pounds to tons?

OpenStudy (anonymous):

What T? Tons

OpenStudy (unklerhaukus):

\[1 [\text{Ton}]=200[\text{lb}]\]\[\frac{1}{200} \left[\frac{\text{Ton}}{\text{lb}}\right]=1\] \[23780 [\text{lb}]=23780 [\text{lb}]\times\frac{1}{200} \left[\frac{\text{Ton}}{\text{lb}}\right]\]\[23780 [\text{lb}]=\frac{23780 [\text{lb}]}{200} \left[\frac{\text{Ton}}{\text{lb}}\right]\] \[23780 [\text{lb}]=\frac{23780 }{200}\left[{\text{Ton}}{}\right]\]